OpenStudy (anonymous):

FAN AND MEDAL Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3

pooja195 (pooja195):

\[\huge~y-y_1=m(x-x_1)\]

pooja195 (pooja195):

m=slope (x1,y1)

OpenStudy (theraggedydoctor):

Btw.. It's *Stephen

OpenStudy (anonymous):

still confused

pooja195 (pooja195):

@StephanHawking Please stop spamming other peoples post. As it is against the CoC http://openstudy.com/code-of-conduct

pooja195 (pooja195):

@kyle55 what part are you confused about?

OpenStudy (anonymous):

So since i have the slope of -3 and the one point what do i do first

pooja195 (pooja195):

\[y-y1=-3(x-x1)\]

OpenStudy (anonymous):

so that would be y- 6 = -3(x + 1) ?

pooja195 (pooja195):

yep :)

OpenStudy (anonymous):

Then from there distribute?

OpenStudy (anonymous):

no

pooja195 (pooja195):

Well it said the point slope form

pooja195 (pooja195):

So i dont think so :/

OpenStudy (anonymous):

oh so thats it

pooja195 (pooja195):

yes :)

OpenStudy (anonymous):

tyvm

OpenStudy (anonymous):

can u check my next question?

pooja195 (pooja195):

sure :)

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1).

OpenStudy (anonymous):

what do i do first? same equation?

pooja195 (pooja195):

First we need to find the slope do you know how?

OpenStudy (anonymous):

y1-y2 over x1-x2

pooja195 (pooja195):

Yes so find the slope first :)

OpenStudy (anonymous):

1 over -1?

OpenStudy (anonymous):

or -1

pooja195 (pooja195):

yes

pooja195 (pooja195):

Now we plug it into the same equation lets use the point (-3,2) and the slope as -1

OpenStudy (anonymous):

so y - 2 = -1(x + 3

pooja195 (pooja195):

yep

OpenStudy (anonymous):

a. y + 3 = −5(x − 2) b. y − 2 = −5(x + 3) c. y + 3 = −one fifth(x − 2) d. y − 2 = −one fifth(x + 3)

OpenStudy (anonymous):

those were the answers

pooja195 (pooja195):

ooo our slope must be wrong :/ let me double check

pooja195 (pooja195):

it would be -1/5

OpenStudy (anonymous):

so d

pooja195 (pooja195):

yep :) The rest of the equation remains the same

OpenStudy (anonymous):

got more time?

pooja195 (pooja195):

I can help with more but repost it :)

pooja195 (pooja195):

On a diffrent thread :P

OpenStudy (anonymous):

ill post my last 2 questions

pooja195 (pooja195):

ok :)